# Learning World Models for Interactive Video Generation

### v1

@@ 3.1 Preliminary: Latent Video Diffusion Model
Video diffusion models have emerged as a powerful framework for video generation. We adopt a latent video diffusion model blattmann2023stable that operates in a compressed latent space rather than pixel space for computational efficiency. Specifically, given an input video sequence $\boldsymbol{x}\in\mathbb{R}^{L\times H\times W\times 3}$, where $L$ is the number of frames, $H$ and $W$ are the height and width of each frame respectively, we first encode it into a latent representation $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ using a pretrained variational autoencoder (VAE). The forward process gradually adds Gaussian noise to the latent according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$:
@@ 3.2 Interactive Long Video Generation
To enable interactive long video generation conditioned on action sequences, we augment the base diffusion model with two techniques: (1) additional action condition with adaptive layer normalization (AdaLN), and (2) random frame noise for autoregressive modeling.
@@
To enable interactive video generation conditioned on action sequences, we augment the base diffusion model with adaptive layer normalization (AdaLN). Given an action sequence $\boldsymbol{a}\in\mathbb{R}^{L\times A}$ where $A$ is the action dimension, we first embed it into a latent space using a learnable embedding layer: $e_{a}=\text{Embed}(\boldsymbol{a})\in\mathbb{R}^{L\times D_{e}}$ where $D_{e}$ is the embedding dimension. For each normalization layer in the diffusion model, we learn action-dependent scale and shift parameters through linear projections: $\gamma_{a}=W_{\gamma}e_{a}+b_{\gamma}\in\mathbb{R}^{L\times D_{h}},\beta_{a}=W_{\beta}e_{a}+b_{\beta}\in\mathbb{R}^{L\times D_{h}}$, where $D_{h}$ matches the hidden dimension of the feature maps. The AdaLN operation is then defined as $\text{AdaLN}(h)=\gamma_{a}\odot\text{LayerNorm}(h)+\beta_{a}$, where $h\in\mathbb{R}^{L\times D_{h}}$ represents the intermediate feature maps and $\odot$ denotes dot production. These AdaLN layers are integrated into each DiT block of the diffusion model, allowing the network to modulate its feature representations based on the input actions while maintaining the core diffusion architecture.
Auto-regressive Video Generation
@@
Specifically, during training, we randomly add noise to each frame in the entire input video sequence according to the diffusion schedule: $z^{i}_{t}=\sqrt{\bar{\alpha}_{t}}z^{i}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},\epsilon^{i}\sim\mathcal{N}(0,\mathbf{I})$, where $z^{i}_{t}$ represents the noised latent of the $i$-th frame. This forces the model to be robust to noise in the conditioning frames and prevents it from relying too heavily on the context. With above two techniques, the training objective for action-conditioned auto-regressive video models become:
@@ 3.3 Retrieval Augmented Video World Model with Global State
where $\tilde{t}$ is a concatenation of $[t^{\prime}]$ and $[t]$, with $t^{\prime}<t$ and $t^{\prime},t\in[T]$.
@@ 3.4 Long-context Extension Baselines
We extend the temporal context window using YaRN peng2023yarn modification for RoPE in temporal attention. RoPE encodes relative positions via complex-valued rotations, where the inner product between query $\mathbf{q}_{m}$ and key $\mathbf{k}_{n}$ depends on relative distance $(m-n)$. YaRN extends the context window by applying a frequency transformation to the rotary position embeddings. This transformation scales the rotation angles in a way that preserves the relative positioning information while allowing the model to handle longer video sequences, after small-scale fine-tuning on longer video clips.
@@ 4.3 World Coherence Results
We investigate the spatiotemporal consistency of internal world models by evaluating the predicted videos given initial frames and action sequences. As visualized in Fig. 4, our VRAG provides an effective approach to enhance the model’s ability to leverage historical context for improving world coherence. Fig. 3 shows more visual comparison of VRAG with ground truth videos. We evaluate the world coherence of different methods using multiple metrics. Figure 5 shows the SSIM scores over time, while Table 1 presents a comprehensive comparison across all metrics. Our VRAG method achieves the best performance across all metrics, demonstrating its superior ability to maintain world coherence in generated videos. Our experimental results demonstrate that expanding the window size from 10 to 20 frames in the baseline DF model improves world coherence, indicating that longer context windows enhance consistency. However, further context extension using YaRN shows no improvement over the vanilla DF model. This suggests that YaRN’s context extension capabilities, while effective in language models, do not transfer effectively to video generation for maintaining world coherence. Similarly, the History Buffer method fails to effectively utilize historical frames for spatiotemporal consistency without explicit in-context training. These findings from both YaRN and History Buffer approaches reveal that video diffusion models at the current scale possess limited in-context learning capabilities, preventing them from effectively leveraging historical frames for maintaining long-term consistency. The Neural Memory method performs poorly due to its instability in model training.
@@
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
DF (window 10) 0.455 16.161 0.509
DF (window 20) 0.466 16.643 0.538
YaRN 0.462 16.567 0.532
History Buffer 0.459 16.922 0.543
VRAG 0.506 17.097 0.506
@@ 4.4 Compounding Error Results
We evaluate the compounding error in long video generation across different methods using the SSIM metric. As shown in Figure 7 and Table 2, our VRAG method achieves superior performance with an SSIM score of 0.349, demonstrating better structural similarity preservation compared to baseline methods. Increasing the window size in DF from 10 to 20 frames improves SSIM, indicating that longer context helps mitigate compounding errors. However, this improvement is still inferior to VRAG’s performance, suggesting that our retrieval-augmented approach provides more effective long-term consistency. As visualized in Fig. 6, our VRAG method generates more coherent and consistent frames over long sequences, while other methods exhibit noticeable artifacts and inconsistencies. The History Buffer method performs poorly, with an SSIM score of 0.188, indicating that naive historical frame retrieval without effective in-context training fails to maintain long-term consistency. Given its limited performance in the world coherence experiments (Section 4.3), we exclude the Neural Memory method from visualization in this longer video prediction visualization.
@@
4.5 Ablation: Memory and Training of VRAG
We ablate the key designs for VRAG methods, including the memory and training components. The ablation results are shown in Figure 8 and Table 3. We compare the full VRAG model with two variants: (1) VRAG without the memory component, which only uses additional global state conditioning, and (2) VRAG without the training component, i.e., vanilla model with retrieval augmented generation for in-context learning at inference. The ablation study is conducted on the world coherence evaluation dataset.
@@ Appendix A Baseline Method Details
To extend the temporal context window of our video generation model, we apply the YaRN [59] modification for ROPE in temporal attention module for improved extrapolation. RoPE encodes relative position via complex-valued rotations, such that the inner product between the $m$-th query $\mathbf{q}_{m}$ and $n$-th key $\mathbf{k}_{n}$ depends only on the relative distance $(m-n)$:
@@
where $\sigma(\cdot)$ is an element-wise nonlinear activation function (e.g., ELU$(\cdot)+1$) and $\mathbf{n}_{s-1}$ is a normalization vector (accumulated up to segment $s-1$).
@@ C.1 Analysis of Compounding Error Evaluation Metrics
To ensure a fair comparison, we evaluate YaRN against a baseline that directly extrapolates the vanilla model’s window size from 20 during training to 40 at inference, to match the inference window length as YaRN in our experiments as Sec. 4. Evaluation of quantitative metrics LPIPS, SSIM and PSNR shown in Figures 1011, and 12 indicates that, YaRN maintains lower compounding error for long video generation (1100 frames). This demonstrates YaRN’s effectiveness in extending the context window of diffusion video models to 40 frames after minimal fine-tuning. Vanilla extension of context length on DF models performs poorly due to out-of-distribution window size at inference.

### v2

@@ Abstract
Project page: https://sites.google.com/view/vrag.
@@ 3.1 Preliminary: Latent Video Diffusion Model
Video diffusion models have emerged as a powerful framework for video generation. We adopt a latent video diffusion model blattmann2023stable that operates in a compressed latent space rather than pixel space for computational efficiency. Specifically, given an input video sequence $\boldsymbol{x}\in\mathbb{R}^{L\times H\times W\times 3}$, we first encode it into a latent representation $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ using a pretrained variational autoencoder (VAE). The forward process gradually adds Gaussian noise to the latent according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$:
@@ 3.2 Interactive Long Video Generation
To enable interactive long video generation conditioned on action sequences, we augment the base diffusion model with two techniques: (1) additional action condition with adaptive layer normalization (AdaLN), and (2) random frame noise for autoregressive modeling, as shown in diagram Fig. 2.
@@
To enable interactive video generation conditioned on action sequences, we augment the base diffusion model with adaptive layer normalization (AdaLN). Given an action sequence $\boldsymbol{a}\in\mathbb{R}^{L\times A}$ where $A$ is the action dimension, we first embed it into a latent space using a learnable embedding layer: $e_{a}=\text{Embed}(\boldsymbol{a})\in\mathbb{R}^{L\times D_{e}}$ where $D_{e}$ is the embedding dimension. For each normalization layer in the diffusion model, we learn action-dependent scale and shift parameters through linear projections: $\gamma_{a}=e_{a}W_{\gamma}+b_{\gamma}\in\mathbb{R}^{L\times D_{h}},\beta_{a}=e_{a}W_{\beta}+b_{\beta}\in\mathbb{R}^{L\times D_{h}}$, where $D_{h}$ matches the hidden dimension of the feature maps. We have $\text{AdaLN}(h)=\gamma_{a}\odot\text{LayerNorm}(h)+\beta_{a}$, where $h\in\mathbb{R}^{L\times D_{h}}$ represents the intermediate feature maps and $\odot$ denotes dot production.
Autoregressive Video Generation
@@
Specifically, during training, we randomly add noise to each frame in the entire input video sequence according to the diffusion schedule: $z^{i}_{t}=\sqrt{\bar{\alpha}_{t}}z^{i}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},\epsilon^{i}\sim\mathcal{N}(0,\mathbf{I})$, where $z^{i}_{t}$ represents the noised latent of the $i$-th frame. This forces the model to be robust to noise in the conditioning frames and prevents it from relying too heavily on the context. With above two techniques, the training objective for action-conditioned autoregressive video models become:
@@ 3.3 Retrieval Augmented Video World Model with Global State
where $\tilde{t}$ is a concatenation of $t^{\prime}]$ and $[t]$, with $t^{\prime}<t$ and $t^{\prime},t\in[T]$.
@@ 3.4 Long-context Extension Baselines
We extend the temporal context window using YaRN peng2023yarn modification for RoPE in temporal attention. RoPE encodes relative positions via complex-valued rotations, where the inner product between query $\mathbf{q}_{m}$ and key $\mathbf{k}_{n}$ depends on relative distance (mn)(m-n). YaRN extends the context window by applying a frequency transformation to the rotary position embeddings. This transformation scales the rotation angles in a way that preserves the relative positioning information while allowing the model to handle longer video sequences, after small-scale fine-tuning on longer video clips.
@@
Frame Pack
As another baseline, we follow the Frame Pack (zhang2025packinginputframecontext, ) to compress historical frames as context. Three input compression kernels with different kernel sizes-(2, 4, 4), (4, 8, 8), and (8, 16, 16)-are employed to condense the historical frames into a fixed-length context. This approach essentially achieves frame compression through importance sampling with recency bias, which enables a larger field of view while maintaining lower computational costs. However, the prioritization of most recent frames can be suboptimal in many cases for long video generation especially when considering the memory issue. Our VRAG based on frame relevance provides theoretically better historical information retrieval. Moreover, our method is actually orthogonal to the frame compression technique in Frame Pack. We leave the combined methods as future work.
@@ 4.3 World Coherence Results
We investigate the spatiotemporal consistency of internal world models by evaluating the predicted videos given initial frames and action sequences. As visualized in Fig. 4, our VRAG provides an effective approach to enhance the model’s ability to leverage historical context for improving world coherence. Fig. 3 shows more visual comparison of VRAG with ground truth videos. We evaluate the world coherence of different methods using multiple metrics. Fig. 5 shows the SSIM scores over time, while Tab. 1 presents a comprehensive comparison across all metrics. Our VRAG method achieves the best performance across all metrics, demonstrating its superior ability to maintain world coherence in generated videos. Our experimental results demonstrate that expanding the window size from 10 to 20 frames in the baseline DF model improves world coherence, indicating that longer context windows enhance consistency. However, further context extension using YaRN shows no improvement over the vanilla DF model. This suggests that YaRN’s context extension capabilities, while effective in language models, do not transfer effectively to video generation for maintaining world coherence. Similarly, the History Buffer method fails to effectively utilize historical frames for spatiotemporal consistency without explicit in-context training. These findings from both YaRN and History Buffer approaches reveal that video diffusion models at the current scale possess limited in-context learning capabilities, preventing them from effectively leveraging historical frames for maintaining long-term consistency. The Neural Memory method performs poorly due to its instability in model training.
@@
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DF (window 10) 0.455 16.161 0.509 DF (window 20) 0.466 16.643 0.538 YaRN 0.462 16.567 0.532 History Buffer 0.459 16.922 0.543 Frame Pack 0.421 16.372 0.574 VRAG 0.506 17.097 0.506
@@ 4.4 Compounding Error Results
We evaluate the compounding error in long video generation across different methods using the SSIM metric. As shown in Fig. 7 and Tab. 2, our VRAG method achieves superior performance with an SSIM score of 0.349, demonstrating better structural similarity preservation compared to baseline methods. Increasing the window size in DF from 10 to 20 frames improves SSIM, indicating that longer context helps mitigate compounding errors. However, this improvement is still inferior to VRAG’s performance, suggesting that our retrieval-augmented approach provides more effective long-term consistency. As visualized in Fig. 6, our VRAG method generates more coherent and consistent frames over long sequences, while other methods exhibit noticeable artifacts and inconsistencies. The History Buffer method performs poorly, with an SSIM score of 0.188, indicating that naive historical frame retrieval without effective in-context training fails to maintain long-term consistency. Given its limited performance in the world coherence experiments (Sec. 4.3), we exclude the Neural Memory method from visualization in this longer video prediction visualization.
@@
4.5 VBench Evaluation
We evaluate the long video generation with five Video Quality metrics in VBench (10657096, ) for generated videos in Sec. 4.4. The evaluation results on VBench (higher is better) are shown in the Tab. 3. As demonstrated in the results, our method outperforms all other baselines across all metrics in both temporal quality and video frame quality. The Neural Memory baseline has Aesthetic Quality 0.343 and Imaging Quality score 0.3597 respectively, which are significantly lower than other baselines, therefore not listed here. Our VRAG shows better temporal consistency compared with all baseline methods, and the high video frame quality indicates the results are not over-smoothed.
4.6 Extension: Real World Setting
Figure 8: Visualized video frames on RealEstate10K dataset. Red blocks indicate the ground-truth frames.
Method Background Consistency Temporal Flickering Motion Smoothness Aesthetic Quality Imaging Quality DF20 0.9668 0.9485 0.9582 0.5272 0.6058 YaRN 0.9686 0.9401 0.9523 0.5252 0.6323 History Buffer 0.9664 0.9475 0.9579 0.5167 0.6253 VRAG 0.9686 0.9511 0.9603 0.5295 0.6444
Table 3: Evaluation results on five Video Quality metrics in VBench.
Metric DFoT VRAG SSIM $\uparrow$ 0.4436 0.9116 PSNR $\uparrow$ 13.03 32.21 LPIPS $\downarrow$ 0.4469 0.1146 FVD $\downarrow$ 337.5 221
Table 4: Quantitative comparison on RealEstate10K dataset.
We conduct additional experiments in real-world setting beyond Minecraft simulation to show generalization of our approach. Specifically, following the experimental setup of Diffusion Forcing Transformer (DFoT) (song2025historyguidedvideodiffusion, ), our VRAG model is initialized from pre-trained DFoT and finetuned on the RealEstate10K dataset (realestate, ) with additionally retrieved historical context, as described in Sec. 3.3.
After fine-tuning for just 2 epochs (10% of the original training steps), our method significantly outperforms the DFoT baseline in terms of memorization capability. The visualized frames are presented in Fig. 8 and quantitative results are summarized in Tab. 4. The results effectively demonstrate the generalization of our approach beyond Minecraft for solving the memory issue in long video prediction.
4.7 Ablation: Memory and Training of VRAG
@@
We ablate the key designs for VRAG methods, including the memory and training components. The ablation results are shown in Fig. 9 and Tab. 5. We compare the full VRAG model with two variants: (1) VRAG without the memory component, which only uses additional global state conditioning, and (2) VRAG without the training component, i.e., vanilla model with retrieval augmented generation for in-context learning at inference. The ablation study is conducted on the world coherence evaluation dataset.
@@ Appendix A Baseline Method Details
To extend the temporal context window of our video generation model, we apply the YaRN [59] modification for ROPE in temporal attention module for improved extrapolation. RoPE encodes relative position via complex-valued rotations, such that the inner product between the $m$-th query $\mathbf{q}_{m}$ and $n$-th key $\mathbf{k}_{n}$ depends only on the relative distance (mn)(m-n):
@@
where $\sigma(\cdot)$ is an element-wise nonlinear activation function (e.g., ELU()+1($\cdot$)+1) and $\mathbf{n}_{s-1}$ is a normalization vector (accumulated up to segment s1s-1).
@@ C.1 Analysis of Compounding Error Evaluation Metrics
To ensure a fair comparison, we evaluate YaRN against a baseline that directly extrapolates the vanilla model’s window size from 20 during training to 40 at inference, to match the inference window length as YaRN in our experiments as Sec. 4. Evaluation of quantitative metrics LPIPS, SSIM and PSNR shown in Figures 1112, and 13 indicates that, YaRN maintains lower compounding error for long video generation (1100 frames). This demonstrates YaRN’s effectiveness in extending the context window of diffusion video models to 40 frames after minimal fine-tuning. Vanilla extension of context length on DF models performs poorly due to out-of-distribution window size at inference.
@@ C.3 More Discussions on Main Results
C.4 Predicted Global State
In the paper, our main experiments are conducted with the access to the ground-truth global state as conditions during training and inference. However, the practical usage may require the global state to be also predicted based on historical states and actions. To ablate this effect, we trained a pose (global state) prediction model that takes the current frame and action as inputs and outputs the predicted pose change. The next post can be derived by adding the predicted pose change to the current pose. Its architecture consists of only a few convolutional layers and fully connected layers, with a very small inference time overhead. At evaluation, we apply this trained predictor to predict the global state at next step, and generate videos based on the predicted global state. Following the same setting as in Sec. 4.3, the experimental results (for 300 frames prediction) are summarized in Table 6.
[…]
DF20 0.466 16.643 0.538
VRAG (predicted pose) 0.500 17.116 0.506
VRAG 0.506 17.097 0.506
Table 6: Ablation study of replacing the ground-truth global state with predicted ones by a trained pose predictor.
As shown in the table, the evaluation results are nearly identical with or without the pose prediction, since the pose prediction is a relatively simple task compared with the video generation. This proves the feasibility of using the predicted global state without significant video performance degradation.
C.5 Memory and Time Overhead
We also compare the memory usage and inference time of VRAG against several baselines: diffusion forcing with 10 and 20 context frames, and YaRN with 40 context frames.
Method DF10 DF20 YaRN VRAG
Context length (Frame) 10 20 40 20
Memory usage (MB) 4420 4448 4543 4452
Inference Time (min) 9 12 23 12
Table 7: Memory usage and inference-time of different methods
As demonstrated in the Table 7, VRAG’s GPU memory usage and inference time overhead are nearly identical to DF20. The inference time is derived for autoregressive generation over 600 frames. Meanwhile, the computation for the retrieval operations can be entirely performed on the CPU, with its memory footprint being only num_frame $\times$ action_dim $\times$ 4 Bytes = 9.4 KB in our experiments, which is almost negligible.
In summary, VRAG incurs almost no additional inference-time overhead compared to standard diffusion forcing. The memory and computational cost introduced by the retrieval mechanism are negligible, as it only involves similarity calculations between a set of vectors.